1. Calculate the pH and pOH of a solution with hydronium ion concentration of 2.5×10-4 M.


2. The pH of a solution is 12.5. What are the hydronium ion and hydroxide ion concentrations in this solution?


1
Expert's answer
2021-06-16T05:32:04-0400

1.) pH = - Log (H3O+)

= - log (2.5 × 10-4)

pH = 3.6

pH + pOH = 14

pOH = 14 - 3.6

pOH = 11.42

2). pH= 12.5

Therefore,

pOH= 14- pH= 14- 12.5= 1.5

Concentration of [OH-]= 10^ -(1.5) =3.163 x10^-2

We Know Kw= [H+] [OH-]= 10^-14 (at25°C)

Hence,

[H+]= Kw/[OH-]= (10^–14)÷(3x10^-2) = 3.162 x 10^-13

[H+]= 3.163 x 10^-2 mol / litre

[OH-] = 3.162 x 10^ -13 mol/ litre